OpenStudy (anonymous):

Suppose all the dimensions of a cube are multiplied by 4. a. what happens to its surface area? b. what happens to its volume?

OpenStudy (anonymous):

 so, SA = 6(6x)^2 = 6*36x^2 = 36*6x^2 the surface area is 36 times the original for volume, V = (6x)^3 = 6*6*6*x^3 = 216*x^3 the volume is 216 times the original volume.